OpenStudy (anonymous):

A disk rotates about its central axis starting from rest and accelerates with constant angular acceleration. At one time it is rotating at 7.8 rev/s; 75 revolutions later, its angular speed is 15 rev/s. Calculate (a) the angular acceleration (rev/s2), (b) the time required to complete the 75 revolutions, (c) the time required to reach the 7.8 rev/s angular speed, and (d) the number of revolutions from rest until the time the disk reaches the 7.8 rev/s angular speed.

OpenStudy (anonymous):

I got in a.) is 1.0944 m/s^2 b.) 6.5785s I got lost in c.. help pls and d maybe

rvc (rvc):

for disc the moment of inertia is : I =MR^2/2

OpenStudy (tylerd):

7.8+at=15rev/s 7.8t+1/2at^2=75 rev

OpenStudy (tylerd):

for time required to complete 75 revolutions we start from rest so its just 1/2at^2=75

OpenStudy (tylerd):

for time required to reach 7.8rev/s at=7.8

OpenStudy (tylerd):

for number of rev from rest you take the time from the last part, at=7.8 and plug it into 1/2at^2

OpenStudy (ac3):

did you get it?